The measure of an angle and its complement are in the ratio 4:11. what is the measure of the two angles? I tried 4x and 11x so far but lost my way of getting to the answer :(
Answer by stanbon(75887)

The measure of an angle and its complement are in the ratio 4:11. what is the measure of the two angles?
---
Let the angle be "x".
Its complement is "90-x"
----
Equation:
x/(90-x) = 4/11
11x = 4(90-x)
11x = 360-4x
15x = 360
----
x = 24 degrees
---
90-x = 66 degrees
===================
Cheers,
Stan H.
=============